OpenStudy (anonymous):

In Jessica's art class, the ratio of girls to boys is 3:2. If there are 25 students in the class, how many more girls are there than boys?

OpenStudy (ladiesman217):

5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

3:2 = Girls:boys girls = 3x 3x + 2x = 25 boys = 2x 5x = 25 x = 5 Girls = 3x = 3(5) = 15 Boys = 2x = 2(5) = 10 15 - 10 = 5

OpenStudy (anonymous):

perfect!!!